OpenStudy (anonymous):

A graduated cylinder with a radius of 3 cm contains water to a height of 9 cm. Yann adds 15 steel balls, each with a diameter of 3 cm. What is the total volume of the contents in the cylinder after the balls are added? Use 3.14 to approximate pi and express your answer in hundredths. This should be 268.47 right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how did you get 268.47?

OpenStudy (anonymous):

I calculated the volume of the cylinder and then the volume of the marbles and added them.

OpenStudy (anonymous):

Oh I forgot to multiply by 15!

jimthompson5910 (jim_thompson5910):

what do you get this time?

OpenStudy (anonymous):

466.29

jimthompson5910 (jim_thompson5910):

what is the volume of the cylinder itself (without any objects inside it)?

OpenStudy (anonymous):

Should be about 254.34

OpenStudy (anonymous):

My answers are correct right?

jimthompson5910 (jim_thompson5910):

volume of cylinder = 254.34 yes, how about the volume of 1 ball?

OpenStudy (anonymous):

About 14.13

jimthompson5910 (jim_thompson5910):

ok, so 15 of them would be ???

OpenStudy (anonymous):

211.95

jimthompson5910 (jim_thompson5910):

so (volume of cylinder) - 15*(volume of 1 ball) = 254.34 - 15*14.13 (volume of cylinder) - 15*(volume of 1 ball) = 254.34 - 211.95 (volume of cylinder) - 15*(volume of 1 ball) = 42.39

OpenStudy (anonymous):

But the question says "What is the total volume of the contents in the cylinder after the balls are ADDED?" Shouldn't I add the volume of the cylinder and the volumes of 15 balls and get 466.29?

jimthompson5910 (jim_thompson5910):

hmm I may be thinking "what is the volume of the water" one sec

jimthompson5910 (jim_thompson5910):

Yeah I'm still thinking it's basically asking "what is the volume of the leftover space in the cylinder after the contents are added?"

OpenStudy (anonymous):

I'll let you know what the answer was

OpenStudy (anonymous):

I was right :) but thanks anyway

jimthompson5910 (jim_thompson5910):

hmm strange question, sorry about that

OpenStudy (anonymous):

It's fine